# biblio.bib Metaheuristics and Large Language Models Join Forces: Towards an Integrated Optimization Approach

### v1

@@ 3.1 Multi-Hop Influence Maximization in Social Networks
For an intuitive explanation of -DSP, see the toy example in Figure 1.
4 Integration of LLM Output into a Metaheuristic
Figure 2 depicts the framework of our proposed MH+LLM integration, comprising three automatic sequential steps:
@@ 4 Integration of LLM Output into a Metaheuristic
4.1 Prompt Engineering
A prompt is an instruction provided as input to an LLM. The design of a prompt can greatly influence the quality of the LLMs’ responses [wei2022emergent, wan2024efficient]. Also, the specific response can vary depending on the LLM used. From the available prompt design techniques, we have opted for one-shot learning [brown2020language], also sometimes called few(1)-shot learning. According to Chen’s findings [chen2023large], tabular data reasoning can achieve good results with a single example, eliminating the need for additional examples or fine-tuning. This method involves furnishing the LLM with an example to facilitate pattern recognition in the response instructions that are to be requested.
4.1.1 Definition
@@
Hereby, the [PROBLEM] tag contains the description of the -DSP. Moreover, the example graph information is provided in the [EXAMPLE GRAPH] tag. Hereby, the example graph consists of 100 nodes, each characterized by the values of five metrics: in-degree, out-degree, closeness, betweenness, and pagerank. In particular, these values are henceforth denoted by
@@
Finally, the [RULES ANSWERING] tag specifies the details of the request to the LLM, which will be elaborated on in Section 4.1.2.
@@
Specifically, contains values for exactly two parameters, regardless of the utilized LLM. The first, known as temperature, is a value between 0 and 1 that measures the model’s response uncertainty, with lower values indicating a more deterministic output. Since a more deterministic response is desirable when searching for patterns in data, we set the temperature to 0.888When generating text or paraphrasing, it is advisable to increase the temperature. This is because creativity is a welcome characteristic in these scenarios. The second parameter is the maximum number of output tokens, which we have set to a moderate tokens. This choice is based on the prompt design, which consistently yields relevant outputs regardless of the evaluation graph’s size, ensuring that the quality of the results is not compromised by a smaller token limit.
4.1.2 Prompt Structure
Effective prompts are generally those with few language ambiguities. To achieve this, the four unique opening and closing tags mentioned in the previous section provide structure and coherence. We will now clarify the syntactic structure of each of these tags. A complete example of a prompt, along with each tag, can be found in Figure 3.
@@ 4.1.2 Prompt Structure
Problem description. The prompt starts by providing a concise definition of the -DSP utilizing LaTeX notation within the [PROBLEM] tag; see the top right of Figure 3.
@@
[DATA] tag: A (directed) random graph with 100 nodes produced with the Erdös– Rényi model [paul1959random] was chosen as an example graph. The edge probability of the graph was . Subsequently, five before-mentioned metrics were calculated for each of the 100 nodes and incorporated into the prompt in a tabular data format, with rows and columns separated by commas. Each row corresponds to a node ID, while the columns represent the respective metric values for that particular node. The metric values are presented in scientific numerical notation to minimize token usage. The rationale behind this decision is discussed in the context of the empirical results; see Section 5.
@@
Rules Answering. The [RULES ANSWERING] tag is crucial as it ties together all the information provided in the previous tags. In this part of the prompt, an equation is presented to the LLM to calculate the probability of each node of the evaluation graph to form part of an optimal solution. The equation requires 10 parameters: 5 alpha parameters and 5 beta parameters, which will be explained in more detail in Section 4.2. These parameters serve to assign weights to the metrics and correct potential errors. The LLM infers the values of these parameters by analyzing the metrics in the [EVALUATION GRAPH] tag and using the metrics and the solution from the [EXAMPLE GRAPH] tag as a guide.
4.2 LLM Output
@@
Note that this formula introduces non-linearity into the node probabilities by applying the sigmoid function , which enables a more nuanced representation of the probability space.999The sigmoid function has been used for many purposes in neural networks. But also in metaheuristics, for example, for significantly accelerating the convergence of a genetic algorithm [4344687].
4.3 Using LLM Output to Guide a Metaheuristic
@@
4.3.1 The Considered Metaheuristic: A BRKGA
The chosen metaheuristic is a so-called Biased Random Key Genetic Algorithm (BRKGA) designed for solving the -DSP in [9909110].101010BRKGA’s are well-known GA variants, mostly for solving combinatorial optimization problems. In fact, two algorithm variants were proposed in [9909110]: (1) a pure BRKGA variant and (2) an algorithm variant that makes use of a hand-designed deep learning framework for biasing the BRKGA. This will allow us to compare our proposal properly to existing algorithm versions.
@@ 4.3.1 The Considered Metaheuristic: A BRKGA
Function Crossover() generates individuals by means of crossover and stores them in set . The crossover, which involves combining two solutions, serves as the mechanism that enhances the search process, concentrating on transferring the superior traits of parents to their offspring.
@@ 4.3.2 Hybrid algorithm
We hypothesize that with suitable predictions from the LLM, the algorithm can be guided/biased to explore more promising areas of the search space. These areas are believed to contain high-quality solutions that the BRKGA would have been unable to discover on their own without the guidance of these predictions. In other words, using an LLM to discover patterns in metric values (see Section 4.1), rather than relying solely on the out-degree, might enhance the algorithm’s performance.
5 Empirical Evaluation
@@
BRKGA: the pure BRKGA variant already published in [9909110] and described above in Section 4.3.1.
BRKGA+FC: the BRKGA hybridized with a hand-designed GNN called FastCover (FC) that was used to derive the probability values (last term of Eq. (11)) in [9909110].
@@
Note that both BRKGA and BRKGA+FC underwent a general parameter tuning in [9909110] depending on the value of . The well-known irace tool [lopez2016irace] was used for this purpose. In this work, we adopt the corresponding parameter settings of BRKGA for BRKGA+LLM. In this way, we can be sure that any difference in their performance is caused by the guidance of the probabilities computed from the LLM outputs. In any case, a specific tuning of BRKGA+LLM could only further improve its results.
@@
5.1 Experimental Setup
The BRKGA was implemented in C++, whereas the prompt construction process, which entails extracting metrics from graph instances, was conducted using Python 3.11. Regarding the choice of LLMs, we utilized two proprietary language models, GPT-4o and Claude-3-Opus, as well as two open-source models, Command-R+ and Mixtral-8x22b-Instruct-v0.1. We selected these models based on the Chatbot Arena—a platform developed by LMSYS members and UC Berkeley SkyLab researchers—which provides an Arena Leaderboard,111111https://chat.lmsys.org. a community-driven ranking system for LLMs [zheng2023judging].121212Please be aware that our experiments took place between February and May 2024, and the LLMs ranking classification in Chatbot Arena may have changed by the time of reading. Table 1 presents a comprehensive overview of the models, including their ranking in the Chatbot Arena Leaderboard (as of May 2024), corresponding version numbers, licenses, maximum context windows, and crucially, the test environment employed for each model.
5.1.1 Execution Environment
We utilized the OpenRouter API131313https://openrouter.ai. to execute prompts in their corresponding LLMs, except for Claude-3-Opus, which we used through the Anthropic API (see Table 1). Finally, all experiments involving the three BRKGA variants were conducted on a high-performance computing cluster comprising machines powered by Intel Xeon CPU 5670 processors with 12 cores running at 2.933 GHz and a minimum of 32 GB of RAM.
5.1.2 Dataset
Our evaluation is based on two categories of -DSP instances (evaluation graphs). The first consists of rather small, synthetic social network graphs with and nodes, generated using three configuration methods developed by Nettleton [Nettleton2016]. The corresponding graph generator requires four real-valued parameters, whose values are reflected by the instance names.141414The instance names are obtained by a concatenation of the utilized parameter values: examples are 0.4-0.15-0.15-0.3, 0.3-0.0-0.3-0.4, and 0.2-0.0-0.3-0.5. The parameters labeled , , , and , respectively, define communities weights ( and ) and link weights between communities ( and ), . These parameters influence the topology of the network, specifically the total number of connections and the density. The second instance set comprises four real-world social network graph instances obtained from the well-established SNAP (Stanford Network Analysis Project) repository [snapnets]. Moreover, note that the -DSP can be solved in each graph for different values of and . In this work we solved all evaluation graphs with and .
@@ 5.1.2 Dataset
The maximum context window of LLMs is still relatively small.151515The maximum context window of an LLM sets the maximum amount of text it can process simultaneously when generating a response. This constraint determines the scope of contextual information the LLM can draw upon when answering a question or completing a task. The response quality will likely degrade if the input prompt exceeds this limit. Given that our prompt design requires each metric for each node to be equally important, the LLM needs to consider as much context as possible to deliver reasonable and trustworthy results. For instance, the largest evaluation graph we use, soc-wiki-elec, results in an input prompt size of tokens, which is close to the tokens limit of Claude-3-Opus [Anthropic2024-ee], the LLM which offers the currently largest context window.
@@
Table 2 provides a detailed breakdown of the constraints for the largest evaluation graphs considered in this work. Although these limitations currently restrict us to testing with smaller instances, we anticipate that this constraint will soon be alleviated as the maximum context window increases and processing costs decrease (see Section 6 for more information on this).
@@
5.1.3 Repository
5.2 Analysis of LLM Output
This section aims to validate the outputs of the LLMs and their usefulness for utilizing them as guidance within the BRKGA algorithm. The aim is to demonstrate that they are not arbitrary or devoid of significance. We have conducted three sets of experiments to achieve this, aiming at different aspects. Figure 4 shows the custom-designed, three-dimensional experimental framework developed specifically for this evaluation.161616Future research could expand the framework’s dimensions to better justify LLMs’ response quality and integration with MHs and to address unexplored aspects.
@@ 5.2 Analysis of LLM Output
Before starting with the main experimental evaluation, we must choose the most suitable LLM for the considered task. For this purpose, we produced all prompts concerning the six synthetic graphs from the dataset (as described before) for all considered combinations of and . These prompts were fed into the following LLMs: GPT-4o, Claude-3-Opus, Command-R+, and Mixtral-8x22b-Instruct-v0.1. The obtained probabilities were then directly used to produce solutions containing the nodes with the largest probability values. The same was done concerning metric out-degree. That is, solutions were produced that consist of the nodes with the highest out-degree values. The results shown in Table 3 allow us to make the following observations. First, although no LLM always outperforms the other LLMs, Claude-3-Opus shows advantages over the other LLMs, especially for increasing values of . When comparing the results obtained with out-degree (the most popular greedy heuristic for the -DSP) to the results obtained with Claude-3-Opus, we can observe that out-degree seems to work slightly better for , while the opposite is the case for . Based on these results, we use Claude-3-Opus for the remainder of our experiments.
5.2.1 Dimension 1 of the Evaluation Framework: Result Quality
In the first set of experiments, we decided to compare the pure BRKGA approach with BRKGA+LLM in the context of the six synthetic graphs (and for all combinations of and ). Both algorithms were applied 10 times to each case, with a computation time limit of 900 CPU seconds per run. The results, which are shown in Table 4, clearly show that BRKGA+LLM outperforms the pure BRKGA algorithm most of the time. Considerable improvements can be observed in the context of the last graph; see instance 0.2-0.0-0.3-0.5 with nodes and arcs. Only in three cases, the result obtained by BRKGA+LLM is slightly inferior to the one of BRKGA. While these results are promising, it is important to recognize that they are based on relatively small instances.
@@ 5.2.1 Dimension 1 of the Evaluation Framework: Result Quality
In the next set of experiments, we applied the BRKGA+FC [9909110], in addition to BRKGA and BRKGA+LLM, to the four larger real-world social networks. Remember that BRKGA+FC is a hybrid approach that uses a hand-crafted GNN approach for biasing the search process of BRKGA. The results are shown—again for each combination of and —in Table 5. The computation time limit for the three approaches was 900 CPU seconds, as in the previously outlined experiments. Moreover, the numbers in the tables are averages over 10 algorithm runs. The results show that BRKGA+LLM outperforms both approaches, with higher margins than those observed in the context of smaller synthetic networks. This holds especially for a growing value of . This is interesting as the prompts only contained an example solution for . This indicates the LLM’s ability to uncover meaningful patterns in the example graph, respectively, in the solution provided in the prompts.
Finally, we aimed to test how meaningful the LLM output really is. For this purpose, we produced two additional variants of BRKGA+LLM: the one called static is obtained by replacing the LLM output with probabilities obtained by random alpha and beta values. The second one, called dynamic, is a similar variant in which the LLM output is replaced with probabilities re-computed at each iteration based on newly determined random alpha and beta values. All three algorithm variants were applied 10 times to each of the four real-world social networks for each combination of and . The results averaged over the 10 runs are shown in Table 6. They clearly indicate that guidance by the LLM output is much more useful than random guidance.
@@
5.2.2 Dimension 2 of the Evaluation Framework: Alternative Techniques for Guiding MHs
To assess the reliability of the LLM output in an alternative way, we decided to compare BRKGA+LLM to an algorithm variant in which the alpha and beta values are obtained by an explicit parameter tuning procedure using the irace tool [lopez2016irace]. This algorithm variant will henceforth be called BRKGA+irace. In particular, for each of the four large social networks—mentioned in Section 5.1.2 and already used, for example, in Table 6—we applied a tuning procedure for obtaining well-working alpha and beta values in the following way. First, for every combination of and a training instance was generated. Second, irace was applied with a budget of algorithm runs, using (as before) a computation time limit of 900 CPU seconds per run. After obtaining the final alpha and beta values from irace for each network, BRKGA+irace was applied under the same conditions as BRKGA+LLM to each of the four problem instances. The results are shown in Table 7. We observe that while BRKGA+irace outperforms BRKGA+LLM in the soc-hamsterster and soc-wiki-elec instances, the opposite is the case for the sign-bitcoinotc instance. In the soc-advogato instance, performance is similar except for , where BRKGA+irace shows better results. For putting these results into perspective, consider that the alpha and beta values of BRKGA+irace were obtained by a specific tuning process that took 250 hours of computing time for each problem instance, while the LLM output (Claude-3-Opus) is obtained in less than a minute without any specific training for our purpose.
We used Pearson’s correlation coefficient [Benesty2009] to identify relationships between the alpha and beta parameters obtained from irace and those from the LLM output (see the rows labeled with in Table 8). A value close to indicates a negative correlation, meaning that when one value increases, the other decreases. A value close to indicates a positive correlation, where both values tend to move together. A value close to means there is no clear relationship between the two series. Our observations are as follows:

### v2

@@ III-A Multi-Hop Influence Maximization in Social Networks
For an intuitive explanation of -DSP, see the toy example in Figure III.
IV Integration of LLM Output into a Metaheuristic
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]figures/overview.pdf An overview of our approach to integrating MHs and LLMs: We employ LLMs to analyze problem instances and uncover hidden patterns. The patterns are then converted into useful information that guides the MH in its search for high-quality solutions.
Figure IV depicts the framework of our proposed MH+LLM integration, comprising three automatic sequential steps:
@@ IV Integration of LLM Output into a Metaheuristic
IV-A Prompt Engineering
A prompt is an instruction provided as input to an LLM. The design of a prompt can greatly influence the quality of the LLMs’ responses [6, 7]. Also, the specific response can vary depending on the LLM used. From the available prompt design techniques, we have opted for one-shot learning [37], also sometimes called few(1)-shot learning. According to Chen’s findings [38], tabular data reasoning can achieve good results with a single example, eliminating the need for additional examples or fine-tuning. This method involves furnishing the LLM with an example to facilitate pattern recognition in the response instructions that are to be requested.
IV-A1 Definition
@@
Hereby, the [PROBLEM] tag contains the description of the -DSP. Moreover, the example graph information is provided in the [EXAMPLE GRAPH] tag. Hereby, the example graph consists of 100 nodes, each characterized by the values of five metrics: in-degree, out-degree, closeness, betweenness, and pagerank.999 We selected these five metrics based on our understanding of the -DSP problem. However, in future work, leveraging the extensive knowledge LLMs possess from academic sources, we plan to use them to guide the selection of metrics for each specific problem. This could represent the extension of our method. In particular, these values are henceforth denoted by
@@
Finally, the [RULES ANSWERING] tag specifies the details of the request to the LLM, which will be elaborated on in Section IV-A2.
@@
Specifically, contains values for exactly two hyperparameters, regardless of the utilized LLM. The first, known as temperature, is a value between 0 and 1 that measures the model’s response uncertainty, with lower values indicating a more deterministic output. While “more deterministic” does not imply that the LLM will generate identical responses to the same prompt every time, it does enhance the stability of the outputs, making it easier to replicate experiments [39]. Therefore, we set the temperature to 0.101010For text generation and paraphrasing tasks, it’s recommended to increase the temperature. This adjustment promotes creativity, which is a desirable characteristic in these contexts. These applications differ from using the LLM as a pattern recognition tool, where consistent and stable outputs are preferable. The second hyperparameter is the maximum number of output tokens, which we have set to a moderate tokens. This choice is based on the prompt design, which consistently yields relevant outputs regardless of the evaluation graph’s size, ensuring that the quality of the results is not compromised by a smaller token limit.
IV-A2 Prompt Structure
Effective prompts are generally those with few language ambiguities. To achieve this, the four unique opening and closing tags mentioned in the previous section provide structure and coherence. We will now clarify the syntactic structure of each of these tags. A complete example of a prompt, along with each tag, can be found in Figure IV-A2.
[p](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]figures/template-prompt.pdf An example of a prompt and the corresponding LLM response. The prompt includes the problem definition, a graph example with node metrics and a high-quality solution, an evaluation graph, and instructions for the LLM for producing the output. Based on the patterns identified in the evaluation graph, the LLM provides the importance of each metric, represented by the set of alpha and beta values.
@@ IV-A2 Prompt Structure
Problem description. The prompt starts by providing a concise definition of the -DSP utilizing LaTeX notation within the [PROBLEM] tag; see the top right of Figure IV-A2.
@@
[DATA] tag: A (directed) random graph with 100 nodes produced with the Erdös–Rényi model [40] was chosen as an example graph. The edge probability of the graph was . Subsequently, five before-mentioned metrics were calculated for each of the 100 nodes and incorporated into the prompt in a tabular data format, with rows and columns separated by commas. Each row corresponds to a node ID, while the columns represent the respective metric values for that particular node. The metric values are presented in scientific numerical notation to minimize token usage. The rationale behind this decision is discussed in the context of the empirical results; see Section V.
@@
Rules Answering. The [RULES ANSWERING] tag is crucial as it ties together all the information provided in the previous tags. In this part of the prompt, an equation is presented to the LLM to calculate the probability of each node of the evaluation graph to form part of an optimal solution. The equation requires 10 parameters: 5 alpha parameters and 5 beta parameters, which will be explained in more detail in Section IV-B. These parameters serve to assign weights to the metrics and correct potential errors. The LLM infers the values of these parameters by analyzing the metrics in the [EVALUATION GRAPH] tag and using the metrics and the solution from the [EXAMPLE GRAPH] tag as a guide.
For optimal prompt construction, we recommend using our OptiPattern tool, which generates the prompt automatically.
IV-B LLM Output
@@
Note that this formula introduces non-linearity into the node probabilities by applying the sigmoid function , which enables a more nuanced representation of the probability space.111111The sigmoid function has been used for many purposes in neural networks. But also in metaheuristics, for example, for significantly accelerating the convergence of a genetic algorithm [41].
IV-C Using LLM Output to Guide a Metaheuristic
@@
IV-C1 The Considered Metaheuristic: A BRKGA
The chosen metaheuristic is a so-called Biased Random Key Genetic Algorithm (BRKGA) designed for solving the -DSP in [19].121212BRKGA’s are well-known GA variants, mostly for solving combinatorial optimization problems. In fact, two algorithm variants were proposed in [19]: (1) a pure BRKGA variant and (2) an algorithm variant that makes use of a hand-designed deep learning framework for biasing the BRKGA. This will allow us to compare our proposal properly to existing algorithm versions.
@@ IV-C1 The Considered Metaheuristic: A BRKGA
A set of individuals are generated by crossover using Crossover() and stored in . The crossover, which involves combining two solutions, serves as the mechanism that enhances the search process, concentrating on transferring the superior traits of parents to their offspring.
@@ IV-C2 Hybrid algorithm
We hypothesize that with suitable predictions from the LLM, the algorithm can be guided/biased to explore more promising areas of the search space. These areas are believed to contain high-quality solutions that the BRKGA would have been unable to discover on their own without the guidance of these predictions. In other words, using an LLM to discover patterns in metric values (see Section IV-A), rather than relying solely on the out-degree, might enhance the algorithm’s performance.
V Empirical Evaluation
@@
Note that both BRKGA and BRKGA+FC underwent a general parameter tuning in [19] depending on the value of . The well-known irace tool [42] was used for this purpose. In this work, we adopt the corresponding parameter settings of BRKGA for BRKGA+LLM. In this way, we can be sure that any difference in their performance is caused by the guidance of the probabilities computed from the LLM outputs. In any case, a specific tuning of BRKGA+LLM could only further improve its results.
@@
V-A Experimental Setup
The BRKGA was implemented in C++, whereas the prompt construction process, which entails extracting metrics from graph instances, was conducted using Python 3.11. Regarding the choice of LLMs, we utilized two proprietary language models, GPT-4o and Claude-3-Opus, as well as two open-source models, Command-R+ and Mixtral-8x22b-Instruct-v0.1. We selected these models based on the Chatbot Arena—a platform developed by LMSYS members and UC Berkeley SkyLab researchers—which provides an Arena Leaderboard,131313https://lmarena.ai. a community-driven ranking system for LLMs [43].141414Please be aware that our experiments took place between February and May 2024, and the LLMs ranking classification in Chatbot Arena may have changed by the time of reading. Table I presents a comprehensive overview of the models, including their ranking in the Chatbot Arena Leaderboard (as of May 2024), corresponding version numbers, licenses, maximum context windows, and crucially, the test environment employed for each model.
V-A1 Execution Environment
We utilized the OpenRouter API151515https://openrouter.ai. to execute prompts in their corresponding LLMs, except for Claude-3-Opus, which we used through the Anthropic API (see Table I). Finally, all experiments involving the three BRKGA variants were conducted on a high-performance computing cluster comprising machines powered by Intel Xeon CPU 5670 processors with 12 cores running at 2.933 GHz and a minimum of 32 GB of RAM.
V-A2 Dataset
Our evaluation is based on two categories of -DSP instances (evaluation graphs). The first consists of rather small, synthetic social network graphs with and nodes, generated using three configuration methods developed by Nettleton [44]. The corresponding graph generator requires four real-valued parameters, whose values are reflected by the instance names.161616The instance names are obtained by a concatenation of the utilized parameter values: examples are 0.4-0.15-0.15-0.3, 0.3-0.0-0.3-0.4, and 0.2-0.0-0.3-0.5. The parameters labeled , , , and , respectively, define communities weights ( and ) and link weights between communities ( and ), . These parameters influence the topology of the network, specifically the total number of connections and the density. The second instance set comprises four real-world social network graph instances obtained from the well-established SNAP (Stanford Network Analysis Project) repository [45]. Moreover, note that the -DSP can be solved in each graph for different values of and . In this work we solved all evaluation graphs with and . All datasets (synthetic and real), prompts, and results can be found in the supplementary material/ folder in the repository: https://github.com/camilochs/optipattern.
@@ V-A2 Dataset
The maximum context window of LLMs is still relatively small.171717The maximum context window of an LLM sets the maximum amount of text it can process simultaneously when generating a response. This constraint determines the scope of contextual information the LLM can draw upon when answering a question or completing a task. The response quality will likely degrade if the input prompt exceeds this limit. Given that our prompt design requires each metric for each node to be equally important, the LLM needs to consider as much context as possible to deliver reasonable and trustworthy results. For instance, the largest evaluation graph we use, soc-wiki-elec, results in an input prompt size of tokens, which is close to the tokens limit of Claude-3-Opus [3], the LLM which offers the currently largest context window.
@@
Table II provides a detailed breakdown of the constraints for the largest evaluation graphs considered in this work. Although these limitations currently restrict us to testing with smaller instances, we anticipate that this constraint will soon be alleviated as the maximum context window increases and processing costs decrease (see Section VI for more information on this).
@@
V-B Analysis of LLM Output
[t!](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]figures/framework.pdf A comprehensive evaluation framework was used to assess the usefulness of integrating MHs with LLMs for solving combinatorial optimization problems (COPs) across the three dimensions shown in the graphic.
This section aims to validate the outputs of the LLMs and their usefulness for utilizing them as guidance within the BRKGA algorithm. The aim is to demonstrate that they are not arbitrary or devoid of significance. We have conducted three sets of experiments to achieve this, aiming at different aspects. Figure V-B shows the custom-designed, three-dimensional experimental framework developed specifically for this evaluation.181818Future research could expand the framework’s dimensions to better justify LLMs’ response quality and integration with MHs and to address unexplored aspects.
@@ V-B Analysis of LLM Output
Before starting with the main experimental evaluation, we must choose the most suitable LLM for the considered task. For this purpose, we produced all prompts concerning the six synthetic graphs from the dataset (as described before) for all considered combinations of and . These prompts were fed into the following LLMs: GPT-4o, Claude-3-Opus, Command-R+, and Mixtral-8x22b-Instruct-v0.1. The obtained probabilities were then directly used to produce solutions containing the nodes with the largest probability values. The same was done concerning metric out-degree. That is, solutions were produced that consist of the nodes with the highest out-degree values. The results shown in Table III allow us to make the following observations. First, although no LLM always outperforms the other LLMs, Claude-3-Opus shows advantages over the other LLMs, especially for increasing values of . When comparing the results obtained with out-degree (the most popular greedy heuristic for the -DSP) to the results obtained with Claude-3-Opus, we can observe that out-degree seems to work slightly better for , while the opposite is the case for . Based on these results, we use Claude-3-Opus for the remainder of our experiments.
V-B1 Dimension 1 of the Evaluation Framework: Result Quality
In the first set of experiments, we decided to compare the pure BRKGA approach with BRKGA+LLM in the context of the six synthetic graphs (and for all combinations of and ). Both algorithms were applied 10 times to each case, with a computation time limit of 900 CPU seconds per run. The results, which are shown in Table IV, clearly show that BRKGA+LLM outperforms the pure BRKGA algorithm most of the time. Considerable improvements can be observed in the context of the last graph; see instance 0.2-0.0-0.3-0.5 with nodes and arcs. Only in three cases, the result obtained by BRKGA+LLM is slightly inferior to the one of BRKGA. While these results are promising, it is important to recognize that they are based on relatively small instances.
@@ V-B1 Dimension 1 of the Evaluation Framework: Result Quality
In the next set of experiments, we applied the BRKGA+FC [19], in addition to BRKGA and BRKGA+LLM, to the four larger real-world social networks. Remember that BRKGA+FC is a hybrid approach that uses a hand-crafted GNN approach for biasing the search process of BRKGA. The results are shown—again for each combination of and —in Table V. The computation time limit for the three approaches was 900 CPU seconds, as in the previously outlined experiments. Moreover, the numbers in the tables are averages over 10 algorithm runs. The results show that BRKGA+LLM outperforms both approaches, with higher margins than those observed in the context of smaller synthetic networks. This holds especially for a growing value of . This is interesting as the prompts only contained an example solution for . This indicates the LLM’s ability to uncover meaningful patterns in the example graph, respectively, in the solution provided in the prompts.
Finally, we aimed to test how meaningful the LLM output really is. For this purpose, we produced two additional variants of BRKGA+LLM: the one called static is obtained by replacing the LLM output with probabilities obtained by random alpha and beta values. The second one, called dynamic, is a similar variant in which the LLM output is replaced with probabilities re-computed at each iteration based on newly determined random alpha and beta values. All three algorithm variants were applied 10 times to each of the four real-world social networks for each combination of and . The results averaged over the 10 runs are shown in Table VI. They clearly indicate that guidance by the LLM output is much more useful than random guidance.
@@
V-B2 Dimension 2 of the Evaluation Framework: Alternative Techniques for Guiding MHs
To assess the reliability of the LLM output in an alternative way, we decided to compare BRKGA+LLM to an algorithm variant in which the alpha and beta values are obtained by an explicit parameter tuning procedure using the irace tool [42]. This algorithm variant will henceforth be called BRKGA+irace. In particular, for each of the four large social networks—mentioned in Section V-A2 and already used, for example, in Table VI—we applied a tuning procedure for obtaining well-working alpha and beta values in the following way. First, for every combination of and a training instance was generated. Second, irace was applied with a budget of algorithm runs, using (as before) a computation time limit of 900 CPU seconds per run. After obtaining the final alpha and beta values from irace for each network, BRKGA+irace was applied under the same conditions as BRKGA+LLM to each of the four problem instances. The results are shown in Table VIII. We observe that while BRKGA+irace outperforms BRKGA+LLM in the soc-hamsterster and soc-wiki-elec instances, the opposite is the case for the sign-bitcoinotc instance. In the soc-advogato instance, performance is similar except for , where BRKGA+irace shows better results. For putting these results into perspective, consider that the alpha and beta values of BRKGA+irace were obtained through a specific tuning process that required significant computational time—approximately 59460 minutes (single-thread) or 2696.7 minutes (parallel jobs on SLURM) for the sum of the four problem instances (see Table VII).191919The irace logs can be found in the supplementary material/ folder in the repository. In contrast, the LLM output (Claude-3-Opus) is obtained in just 5.3 minutes via an API call over HTTP. While this is not a fully fair comparison due to the different hardware setups, the substantial difference in runtime highlights the efficiency of the LLM-based approach. In some scenarios, it may be preferable to pay for API access rather than maintain dedicated servers that require significantly more time to achieve comparable results.
| Time (seconds) | |||
| irace | LLM | ||
| Instance | Single-Thread (Serial) | Multiple Jobs in SLURM (Parallel) | Anthropic API (HTTP) |
| soc-wiki-elec | 891900 | 46422 | 119 |
| soc-advogato | 891900 | 40002 | 114 |
| sign-bitcoinotc | 891900 | 37292 | 51 |
| soc-hamsterster | 891900 | 38087 | 36 |
| Total time (minutes) | 59460.0 | 2696.7 | 5.3 |
We used Pearson’s correlation coefficient [46] to identify relationships between the alpha and beta parameters obtained from irace and those from the LLM output (see the rows labeled with in Table IX). A value close to indicates a negative correlation, meaning that when one value increases, the other decreases. A value close to indicates a positive correlation, where both values tend to move together. A value close to means there is no clear relationship between the two series. Our observations are as follows: